Questions 4 and 5 are based on the following information: The daily rate of sales of a product (in units per day) is approximated by the exponential function

S(t) = 1800 + 1500e−0,3t+1,5


with t the number of days it has been on the market.

Question  4

After 20 days the rate of sales of the product is approximately

 

1.      37.

2.      108.

3.      1 817.

4.      1 783.

 


Question   5

After how many days, rounded to a whole number, will the rate of sale be 2 000 units per day?

 

1.      4

2.      5

3.      11

4.      12


1
Expert's answer
2021-07-12T16:10:04-0400

Question 4

"t=20\\
S(t)=S(20)=1800 + 1500e^{−0.3t+1.5}\\
S(20)=1800 + 1500e^{−0.3*20+1.5}\\
S(20)=1800 + 17\\
S(20)=1817"

Option 3 is correct


Question 5

"S(t) = 1800 + 1500e^{−0.3t+1.5}\\
2 000=1800 + 1500e^{−0.3t+1.5}\\
1800+1500e^{-0.3t+1.5}-1800=2000-1800\\
1500e^{-0.3t+1.5}=200\\
e^{-0.3t+1.5}=\frac{2}{15}\\
t=\frac{-\ln \left(\frac{2}{15}\right)+1.5}{0.3}\\
t=11.7=12"

Option 4 is correct